# A Survey of Mamba

### v2

@@ 3.1. Mamba-1: Selective State Space Model with Hardware-aware Algorithms
Conventional SSMs have shown limited effectiveness in modeling text and other information-dense data (Gu and Dao, 2023), impeding their progress in deep learning. In the pursuit of empowering SSMs with Transformers’ modeling capabilities, Gu and Dao (2023) introduces three innovative techniques based on Structured State Space Models, i.e., High-order Polynomial Projection Operator (HiPPO)-based Memory Initialization, Selection Mechanism, and Hardware-aware Computation, as illustrated in Figure 3. These techniques aim to enhance the capabilities of SSMs in long-range linear-time sequence modeling. In particular, the initialization strategy establishes a coherent hidden state matrix, effectively facilitating long-range memory. Then, the Selection Mechanism empowers SSMs to acquire content-aware representations. Lastly, Mamba crafts two hardware-aware computation algorithms, Parallel Associative Scan and Memory Recomputation, to enhance training efficiency.
@@ 3.1.1. HiPPO-based Memory Initialization
Modeling and learning from sequential data represent foundational challenges in contemporary machine learning, forming the bedrock for various tasks, including language modeling, speech recognition, and video processing. A fundamental component for modeling intricate and long-term temporal dependencies lies in memory, encompassing the ability to store and integrate information from preceding time steps (Hu and Qi, 2017). Similar to RNNs, preserving and forgetting the historical hidden states (i.e., the matrix $\mathbf{A}$) plays a critical role in SSMs to achieve satisfying performances. In previous structured state space sequence models (SSMs), there have been suggestions for special initializations, especially in the case of complex-valued models. These special initializations have proven beneficial in various scenarios, including situations with limited data availability. Similarly, Mamba focuses primarily on the initialization of the hidden state matrix $\mathbf{A}$ to capture complex temporal dependencies. This is accomplished through the utilization of the HiPPO theory (Gu et al., 2020) with an innovative scaled Legendre measure (LegS), ensuring a comprehensive consideration of the complete historical context rather than a limited sliding window. To be specific, the HiPPO-LegS assigns uniform weight to all historical data points, which can be expressed as:
@@ 3.2. Mamba-2: State Space Duality
where $\mathbf{M}$ denotes the matrix form of SSMs that uses the sequentially semi-separable representation, and $\mathbf{M}_{ji}=\mathbf{C}^{\text{T}}_{j}\mathbf{A}_{j:i}\mathbf{B}_{i}$. Notably, $\mathbf{C}_{j}$ and $\mathbf{B}_{i}$ represent the selective space state matrices associated with input tokens $\mathbf{x}_{j}$ and $\mathbf{x}_{i}$, respectively. $\mathbf{A}_{j:i}$ denotes the selective matrix of hidden states corresponding to the input tokens ranging from $j$ to $i$. In essence, SSD demonstrates that both the attention mechanism used by Transformers and the linear time-variant system employed in SSM can be seen as semi-separable matrix transformations. Furthermore, Dao and Gu (2024) also proves that the selective SSM is equivalent to a structured linear attention mechanism implemented with a semi-separable masking matrix.
@@ 4. Advancements in Mamba Models
State Space Models and Mamba have been recently explored and become one promising alternative as the foundational model backbone. As shown in Table 1, large-scale Mamba-based models have not only thrived within academic research but have also made significant strides in industry, such as Falcon Mamba 7B and Mistral 7B, demonstrating their efficacy through successful training on GPUs. Despite that, the Mamba architecture still encounters challenges, such as memory loss, generalization to diverse tasks, and inferior capability to capture complex patterns to Transformer-based language models. To overcome these challenges, plenty of efforts have been made to improve the Mamba architecture. Existing research studies primarily concentrate on modifying the block design, scanning mode, and memory management aspects. This section will introduce several vital techniques from these three aspects, and a summary of related studies is presented in Table 2.
@@
Stereo Hierarchical Scan (Chen et al., 2024f; Wang et al., 2024a; Bhirangi et al., 2024; Chen et al., 2024g; Han et al., 2024; Shi et al., 2024a)
Spatiotemporal Scan (Li et al., 2024c; Chen et al., 2024e; Yao et al., 2024; Yang et al., 2024e)
@@ 4.1. Block Design
Figure 5. Representative examples of improved Mamba models based on the perspective of block design: (a). Integration methods combine orthogonal architectural designs (e.g., Transformer) with Mamba, leading to enhanced model performance and increased throughput, all while upholding a manageable memory footprint (Xu et al., 2024a; Pilault et al., 2024); (b). Substitution methods enhance the learning capabilities of standard learning frameworks (e.g., UNet) by integrating the Mamba block as a substitute for their primary layers (Yue and Li, 2024; Liao et al., 2024); (c). Modification methods employ cutting-edge techniques, such as Mix-of-Expert (MoE), to refine the Mamba block (Lieber et al., 2024).
@@ 4.1.1. Integration
Given Mamba’s exceptional ability to capture long-term dynamics, it has been extensively integrated with other models, leveraging their own strengths to deliver a robust framework tailored to specific scenarios. The integration specifically encompasses advanced models like Transformers, Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), Recurrent Neural Networks (RNNs), and Spiking Neural Networks (SNNs). Specific examples are described below.
@@
  • GNN has demonstrated promising potential in capturing neighboring relationships through message-passing mechanisms, where information is propagated over a connection graph through stacked layers. Nonetheless, these models face a significant limitation known as over-smoothing (Chen et al., 2020), particularly when attempting to capture high-order adjacency signals. To tackle such a challenge, Mamba has been employed for graph representation learning (Liu et al., 2024a; Li et al., 2024f; Yang et al., 2024d; Wang et al., 2024e). For example, Graph Mamba (Behrouz and Hashemi, 2024) reformulates graph-structured data into sequential tokens in a particular order and leverages a selective SSM layer within the Mamba block to construct a novel Graph Mamba Network (GMN) architecture, which achieves superior graph representation learning capabilities, particularly in the datasets that requires high-order dependencies between nodes.
  • SNN has been recently proposed as a promising network architecture inspired by the behavior of biological neurons in the brain: transmitting knowledge between neurons through discrete spikes. One of the key advantages of SNNs lies on their potential for low-power implementation, as they can exploit the sparse and event-driven nature of neural activity. Motivated by the energy-efficient implementation of SNNs and SSMs’ superior long-range learning capabilities, pioneering studies have delved into integrating these two methods. For example, SpikeMba (Li et al., 2024b) combines them to handle confidence bias towards prominent objects and to capture enduring dependencies within video sequences. Through extensive evaluations, the authors claim that integrating these two models improves the effectiveness of temporal video grounding tasks, precisely moment retrieval and highlight detection.
@@ 4.1.2. Substitution
  • U-Net. On the one hand, many efforts (Shi et al., 2024b; Wang et al., 2024g, a; Liao et al., 2024) have been made to synergize U-Net with Mamba’s capability in capturing intricate and broad semantics, so as to advance model performance in computer vision tasks. For example, Mamba-UNet (Wang et al., 2024g) utilizes Visual Mamba blocks exclusively to construct a U-Net-like model (i.e., an encoder-decoder model infused with skip connections) for medical image segmentation. Their evaluation demonstrates that Mamba-UNet surpasses several U-Net variations, which can be attributed to the efficacy and efficiency of Mamba blocks in handling long-range patch sequences.
  • Diffusion Model. On the other hand, some endeavors (Fu et al., 2024; Fei et al., 2024; Oshima et al., 2024) have been undertaken to build a novel type of diffusion model, Diffusion State Space Model (DiS), which replace the typical backbone (e.g., CNNs, Attentions, U-Nets) with a state space backbone. Given the remarkable efficiency and efficacy of Mamba blocks in accommodating long-range dependencies, DiS is distinguished by generating longer sequences using diffusion models (Fei et al., 2024). For example, Oshima et al. (2024) propose a Mamba-based diffusion model that substantially decreases memory consumption for long video sequences, while still maintaining competitive performance metrics when compared to Transformer-based models. Moreover, MD-Dose (Fu et al., 2024) and P-Mamba (Ye and Chen, 2024) construct noise predictors using Mamba blocks in the backward process of diffusion models, ultimately generating specific targets for medical image processing.
  • Others. Besides the U-Net and Diffusion Models, there are a few substitutions. For example, Res-VMamba (Chen et al., 2024b) adopts Visual Mamba blocks in a residual learning framework for food category classification. Furthermore, SPMamba (Li and Chen, 2024) adopts the TF-GridNet (Wang et al., 2023a), a recently developed time-frequency model, as its base architecture followed by succeeding the Transformer components with bidirectional Mamba blocks. This adaptation enable the model to encompass a wider scope of contextual information efficiently for the task of speech separation.
@@ 4.1.3. Modification
Apart from Integrating and Substituting methods that directly employ the Mamba block, some other efforts have been made to modify the Mamba block with the aim of enhancing its performance in different scenarios. For example, Jamba (Lieber et al., 2024) borrows the conception of Mix-of-Experts (MoE) (Jacobs et al., 1991; Fedus et al., 2022) to enable their hybrid (Transformer-Mamba) decoder-only model to be pretrained with far less compute and allow flexible objective-specific configurations. Notably, the Jamba model (56B available parameters, 12B active parameters, 4GB KV cache) requires a 32x smaller KV cache compared to a representative Transformer-based language model, LLaMA-2-7B (6.7B available parameters, 12B active parameters, 128GB KV cache), while providing more extensive available and active parameters. This allows Jamba to swallow a context length of 140K on a single A100 GPU (80GB), seven times the length supported by LLaMA-2-70B. In addition to MoE, some studies propose modifying the SSM layer into a K-way structure, which involves processing model inputs using parallel SSM cells, allowing for capturing information and knowledge from multiple perspectives. For example, Sigma (Wan et al., 2024) develops a novel Mamba-based visual encoder that handles multimodal inputs by utilizing parallel SSM layers. UltraLight VM-UNet (Wu et al., 2024) proposes a vision Mamba layer with parallel SSM cells that processes deep features in different channels. To recap, by implementing such modifications (i.e., K-way, MoE), these Mamba-based models gain enhanced learning capabilities, particularly in processing multimodal inputs and fast adapting to multiscale tasks. In addition, a pioneering study, Mamba®, has introduced a novel approach that suggests incorporating registers evenly within the visual input tokens before passing the inputs through the SSM layers. This modification aims to enhance the representation of the sequence direction of image patches, thereby enabling the unidirectional inference paradigm of the Mamba block to be applicable to visual tasks. Despite these successes, the exploration of modifying Mamba blocks remains a promising yet under-explored area.
@@ 4.2.2. Stereo Scan
  • Hierarchical Scan methods involve employing different kernel sizes of scanning to capture the semantic knowledge from global to local or from macro to micro perspectives (Wang et al., 2024a; Chen et al., 2024g; Han et al., 2024; Shi et al., 2024a). For example, a Mamba-in-Mamba hierarchical encoder is proposed by (Chen et al., 2024f) for infrared small target detection, combining inner and outer selective SSM blocks. The inner one is specifically tailored to capture the interplay among visual patches for local pattern extraction. Conversely, the outer block is designed to characterize the relationship between visual sentences to capture global features. HiSS (Bhirangi et al., 2024) divides an input sequence into chunks and models the chunk features hierarchically for continuous sequential prediction. The chunks are first processed by a low-level SSM cell, and the processed features are mapped into an output sequence by a high-level SSM block.
  • Spatiotemporal Scan. Driven by the prevalence of dynamic systems in the real world, there has been a growing interest in spatiotemporal scanning methods to enhance the performance of Mamba block (Yao et al., 2024; Yang et al., 2024e). For instance, VideoMamba (Li et al., 2024c) expands the original 2D scan for images into two 3D scans: spatial-first scanning and temporal-first scanning. Combining these two scanning approaches, VideoMamba demonstrates exceptional efficiency in handling long, high-resolution videos. Additionally, ChangeMamba (Chen et al., 2024e) integrates three spatiotemporal scanning mechanisms (sequential modeling, cross modeling, and parallel modeling) to enable contextual information interaction among multi-temporal features for remote sensing change detection.
@@ 5. Adapting Mamba to Diverse Data
The Mamba architecture represents an extension of selective state space models, which possesses fundamental properties of recurrent models that make it well-suited as general foundation models operating on sequences like text, time series, speech, and more. Meanwhile, recent pioneering studies have extended the utilization of the Mamba architecture beyond sequential data, encompassing domains such as images and graphs, as depicted in Figure 7. These studies aim to harness Mamba’s remarkable capabilities in capturing long-range dependencies while leveraging its efficiency in learning and inference processes. In this section, we therefore aim to investigate the emerging techniques that adapt Mamba to various types of data.
@@
Speech Speech Separation (Abdel-Hamid et al., 2014; Chen et al., 2024h; Li and Chen, 2024; Jiang et al., 2024b)
@@
Non-Sequential Data Image High-Resolution Medical Vision (Yue and Li, 2024; Chen et al., 2024a, e; Lin et al., 2024b; Wang et al., 2024g, g; Ruan and Xiang, 2024; Zhu et al., 2024; Touvron et al., 2021)

### v3

@@ 3.1. Mamba-1: Selective State Space Model with Hardware-aware Algorithms
Conventional SSMs have shown limited effectiveness in modeling text and other information-dense data (Gu and Dao, 2023), impeding their progress in deep learning. In the pursuit of empowering SSMs with Transformers’ modeling capabilities, Gu and Dao (2023) introduce three innovative techniques based on Structured State Space Models, i.e., High-order Polynomial Projection Operator (HiPPO)-based Memory Initialization, Selection Mechanism, and Hardware-aware Computation, as illustrated in Figure 3. These techniques aim to enhance the capabilities of SSMs in long-range linear-time sequence modeling. In particular, the initialization strategy establishes a coherent hidden state matrix, effectively facilitating long-range memory. Then, the Selection Mechanism empowers SSMs to acquire content-aware representations. Lastly, Mamba crafts two hardware-aware computation algorithms, Parallel Associative Scan and Memory Recomputation, to enhance training efficiency.
@@ 3.1.1. HiPPO-based Memory Initialization
Modeling and learning from sequential data represent foundational challenges in contemporary machine learning, forming the bedrock for various tasks, including language modeling, speech recognition, and video processing. A fundamental component for modeling intricate and long-term temporal dependencies lies in memory, encompassing the ability to store and integrate information from preceding time steps (Hu and Qi, 2017). Similar to RNNs, preserving and forgetting the historical hidden states (i.e., the matrix $\mathbf{A}$) play a critical role in SSMs to achieve satisfying performances. In previous structured state space sequence models (SSMs), there have been suggestions for special initializations, especially in the case of complex-valued models. These special initializations have proven beneficial in various scenarios, including situations with limited data availability. Similarly, Mamba focuses primarily on the initialization of the hidden state matrix $\mathbf{A}$ to capture complex temporal dependencies. This is accomplished through the utilization of the HiPPO theory (Gu et al., 2020) with an innovative scaled Legendre measure (LegS), ensuring careful consideration of the complete historical context rather than a limited sliding window. To be specific, the HiPPO-LegS assigns uniform weight to all historical data points, which can be expressed as:
@@ 3.2. Mamba-2: State Space Duality
where $\mathbf{M}$ denotes the matrix form of SSMs that uses the sequentially semi-separable representation, and $\mathbf{M}_{ji}=\mathbf{C}^{\text{T}}_{j}\mathbf{A}_{j:i}\mathbf{B}_{i}$. Notably, $\mathbf{C}_{j}$ and $\mathbf{B}_{i}$ represent the selective space state matrices associated with input tokens $\mathbf{x}_{j}$ and $\mathbf{x}_{i}$, respectively. $\mathbf{A}_{j:i}$ denotes the selective matrix of hidden states corresponding to the input tokens ranging from $j$ to $i$. In essence, SSD demonstrates that both the attention mechanism used by Transformers and the linear time-variant system employed in SSM can be seen as semi-separable matrix transformations. Furthermore, Dao and Gu (2024) also prove that the selective SSM is equivalent to a structured linear attention mechanism implemented with a semi-separable masking matrix.
@@ 4. Advancements in Mamba Models
State Space Models and Mamba have been recently explored and have become one promising alternative as the foundational model backbone. As shown in Table 2, large-scale Mamba-based models have not only thrived within academic research but have also made significant strides in industry, such as Falcon Mamba 7B and Mistral 7B, demonstrating their efficacy through successful training on GPUs. Despite that, the Mamba architecture still encounters challenges, such as memory loss, generalization to diverse tasks, and inferior capability to capture complex patterns to Transformer-based language models. To overcome these challenges, plenty of efforts have been made to improve the Mamba architecture. Existing research studies primarily concentrate on modifying the block design, scanning mode, and memory management aspects. This section will introduce several vital techniques from these three aspects, and a summary of related studies is presented in Table 3.
@@
Stereo Hierarchical Scan (Chen et al., 2024e; Wang et al., 2024a; Bhirangi et al., 2024; Chen et al., 2024f; Han et al., 2024; Shi et al., 2024a)
Spatiotemporal Scan (Li et al., 2024c; Chen et al., 2024d; Yao et al., 2024; Yang et al., 2024e)
@@ 4.1. Block Design
Figure 5. Representative examples of improved Mamba models based on the perspective of block design: (a) Integration methods combine orthogonal architectural designs (e.g., Transformer) with Mamba, leading to enhanced model performance and increased throughput, all while upholding a manageable memory footprint (Xu et al., 2024a; Pilault et al., 2024); (b) Substitution methods enhance the learning capabilities of standard learning frameworks (e.g., UNet) by integrating the Mamba block as a substitute for their primary layers (Yue and Li, 2024; Liao et al., 2024); (c) Modification methods employ cutting-edge techniques, such as Mix-of-Expert (MoE), to refine the Mamba block (Lieber et al., 2024).
@@ 4.1.1. Integration
Given Mamba’s exceptional ability to capture long-term dynamics, it has been extensively integrated with other models, leveraging its strengths to deliver a robust framework tailored to specific scenarios. The integration specifically encompasses advanced models like Transformers, Graph Neural Networks (GNNs), Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), and Spiking Neural Networks (SNNs). Specific examples are described below.
@@
  • GNN has demonstrated promising potential in capturing neighboring relationships through message-passing mechanisms, where information is propagated over a connection graph through stacked layers. Nonetheless, these models face a significant limitation known as over-smoothing (Chen et al., 2020), particularly when attempting to capture high-order adjacency signals. To tackle such a challenge, Mamba has been employed for graph representation learning (Liu et al., 2024a; Li et al., 2024f; Yang et al., 2024d; Wang et al., 2024e). For example, Graph Mamba (Behrouz and Hashemi, 2024) reformulates graph-structured data into sequential tokens in a particular order and leverages a selective SSM layer within the Mamba block to construct a novel Graph Mamba Network (GMN) architecture, which achieves superior graph representation learning capabilities, particularly in the datasets that require high-order dependencies between nodes.
@@
  • SNN has been recently proposed as a promising network architecture inspired by the behavior of biological neurons in the brain: transmitting knowledge between neurons through discrete spikes. One of the key advantages of SNNs lies in their potential for low-power implementation, as they can exploit the sparse and event-driven nature of neural activity. Motivated by the energy-efficient implementation of SNNs and SSMs’ superior long-range learning capabilities, pioneering studies have delved into integrating these two methods. For example, SpikeMba (Li et al., 2024b) combines them to handle confidence bias towards prominent objects and to capture enduring dependencies within video sequences. Through extensive evaluations, the authors claim that integrating these two models improves the effectiveness of temporal video grounding tasks, precisely moment retrieval and highlight detection.
@@ 4.1.2. Substitution
  • U-Net. Many efforts (Shi et al., 2024b; Wang et al., 2024g, a; Liao et al., 2024) have been made to synergize U-Net with Mamba’s capability in capturing intricate and broad semantics so as to advance model performance in computer vision tasks. For example, Mamba-UNet (Wang et al., 2024g) utilizes Visual Mamba blocks exclusively to construct a U-Net-like model (i.e., an encoder-decoder model infused with skip connections) for medical image segmentation. Their evaluation demonstrates that Mamba-UNet surpasses several U-Net variations, which can be attributed to the efficacy and efficiency of Mamba blocks in handling long-range patch sequences.
  • Diffusion Model. Some endeavors (Fu et al., 2024; Fei et al., 2024; Oshima et al., 2024) have been undertaken to build a novel type of diffusion model, Diffusion State Space Model (DiS), which replace the typical backbone (e.g., CNNs, Attentions, U-Nets) with a state space backbone. Given the remarkable efficiency and efficacy of Mamba blocks in accommodating long-range dependencies, DiS is distinguished by generating longer sequences using diffusion models (Fei et al., 2024). For example, Oshima et al. (2024) propose a Mamba-based diffusion model that substantially decreases memory consumption for long video sequences, while still maintaining competitive performance metrics when compared to Transformer-based models. Moreover, MD-Dose (Fu et al., 2024) and P-Mamba (Ye and Chen, 2024) construct noise predictors using Mamba blocks in the backward process of diffusion models, ultimately generating specific targets for medical image processing.
  • Others. Besides the U-Net and Diffusion Models, there are a few substitutions. For example, Res-VMamba (Chen et al., 2024b) adopts Visual Mamba blocks in a residual learning framework for food category classification. Furthermore, SPMamba (Li and Chen, 2024) adopts the TF-GridNet (Wang et al., 2023a), a recently developed time-frequency model, as its base architecture followed by succeeding the Transformer components with bidirectional Mamba blocks. This adaptation enables the model to encompass a wider scope of contextual information efficiently for the task of speech separation.
@@ 4.1.3. Modification
Apart from integration and substitution methods that directly employ the Mamba block, some other efforts have been made to modify the Mamba block with the aim of enhancing its performance in different scenarios. For example, Jamba (Lieber et al., 2024) borrows the conception of Mix-of-Experts (MoE) (Jacobs et al., 1991; Fedus et al., 2022) to enable their hybrid (Transformer-Mamba) decoder-only model to be pretrained with far less compute and allow flexible objective-specific configurations. Notably, the Jamba model (56B available parameters, 12B active parameters, 4GB KV cache) requires a 32x smaller KV cache compared to a representative Transformer-based language model, LLaMA-2-7B (6.7B available parameters, 12B active parameters, 128GB KV cache), while providing more extensive available and active parameters. This allows Jamba to swallow a context length of 140K on a single A100 GPU (80GB), seven times the length supported by LLaMA-2-70B. In addition to MoE, some studies propose modifying the SSM layer into a K-way structure, which involves processing model inputs using parallel SSM cells, allowing for capturing information and knowledge from multiple perspectives. For example, Sigma (Wan et al., 2024) develops a novel Mamba-based visual encoder that handles multimodal inputs by utilizing parallel SSM layers. UltraLight VM-UNet (Wu et al., 2024) proposes a vision Mamba layer with parallel SSM cells that process deep features in different channels. To recap, by implementing such modifications (i.e., K-way, MoE), these Mamba-based models gain enhanced learning capabilities, particularly in processing multimodal inputs and fast adapting to multiscale tasks. In addition, a pioneering study, Mamba®, has introduced a novel approach that suggests incorporating registers evenly within the visual input tokens before passing the inputs through the SSM layers. This modification aims to enhance the representation of the sequence direction of image patches, thereby enabling the unidirectional inference paradigm of the Mamba block to be applicable to visual tasks. Despite these successes, the exploration of modifying Mamba blocks remains a promising yet under-explored area.
@@ 4.2.2. Stereo Scan
  • Hierarchical Scan methods involve employing different kernel sizes of scanning to capture the semantic knowledge from global to local or from macro to micro perspectives (Wang et al., 2024a; Chen et al., 2024f; Han et al., 2024; Shi et al., 2024a). For example, a Mamba-in-Mamba hierarchical encoder is proposed by (Chen et al., 2024e) for infrared small target detection, combining inner and outer selective SSM blocks. The inner one is specifically tailored to capture the interplay among visual patches for local pattern extraction. Conversely, the outer block is designed to characterize the relationship between visual sentences to capture global features. HiSS (Bhirangi et al., 2024) divides an input sequence into chunks and models the chunk features hierarchically for continuous sequential prediction. The chunks are first processed by a low-level SSM cell, and the processed features are mapped into an output sequence by a high-level SSM block.
  • Spatiotemporal Scan. Driven by the prevalence of dynamic systems in the real world, there has been a growing interest in spatiotemporal scanning methods to enhance the performance of Mamba block (Yao et al., 2024; Yang et al., 2024e). For instance, VideoMamba (Li et al., 2024c) expands the original 2D scan for images into two 3D scans: spatial-first scanning and temporal-first scanning. Combining these two scanning approaches, VideoMamba demonstrates exceptional efficiency in handling long, high-resolution videos. Additionally, ChangeMamba (Chen et al., 2024d) integrates three spatiotemporal scanning mechanisms (sequential modeling, cross modeling, and parallel modeling) to enable contextual information interaction among multi-temporal features for remote sensing change detection.
@@ 5. Adapting Mamba to Diverse Data
The Mamba architecture represents an extension of selective state space models, which possesses fundamental properties of recurrent models that make it well-suited as general foundation models operating on sequences like text, time series, speech, and more. Meanwhile, recent pioneering studies have extended the utilization of the Mamba architecture beyond sequential data, encompassing domains such as images and graphs, as depicted in Figure 7. These studies aim to harness Mamba’s remarkable capabilities in capturing long-range dependencies while leveraging its efficiency in learning and inference processes. In this section, we therefore aim to investigate the emerging techniques that adapt Mamba to various types of data. A summary of related studies is illustrated in Table 4.
@@
Speech Speech Separation (Abdel-Hamid et al., 2014; Chen et al., 2024g; Li and Chen, 2024; Jiang et al., 2024b)
@@
Non-Sequential Data Image High-Resolution Medical Vision (Yue and Li, 2024; Chen et al., 2024a, d; Lin et al., 2024b; Wang et al., 2024g, g; Ruan and Xiang, 2024; Zhu et al., 2024; Touvron et al., 2021)